If f(x)=x^2-3x+5, find f(1)-f(4)
Answer by calbreez(9)

f(1)=1^2-3*1+5
=1-3+5
=3
f(4)=4^2-3*4+5
=16-12+5
=9
So f(1)-f(4)=3-9
=-6